OpenStudy (anonymous):

sorry i cant get it right on here but i'm looking for the domain: f(x2) x+2 over x^3-3x^2-4x12 domain is{x|x not equal sign______________}

OpenStudy (anonymous):

Is this it?\[f(x ^{2})=\left(\begin{matrix}x+2 \\ x ^{3}-3x ^{2}-4*12\end{matrix}\right)\]

OpenStudy (anonymous):

almost. hi dryanni...sorry it looks confusing but it's like this: f(x)= x+2 ________________ x^3-3x^2-4x+12

OpenStudy (anonymous):

there it goes again..lol....that line is suppossed to be the "over" sign f(x)= x+2 (these would be the numerator) over sign and then the denominator would be x^3-3x^2-4x+12

OpenStudy (anonymous):

Both the numerator and the denominator are defined for the entire domain of x. The equation is not defined (≠) when \[x^3-3x^2-4x+12\]is equal to 0, because you cannot divide by 0. This is therefore a root problem for a polynomial of the third degree.

OpenStudy (mimi_x3):

lols , next time just do "/" for over xD

OpenStudy (anonymous):

oh goodness...i was looking for the right sign....dumbcow showed me another sign before too...very helpful. i'll get it right next time.

OpenStudy (anonymous):

okay let me try it again f(x) x+2 / x^3-3x^2-4x+12 domain: {x|x and not equal sign then the answer ___________}